A bicyclist rode into the country for 5h. In returning, her speed was 5mi/h faster and the trip took 4h. what was her speed each way?
Answer by ankor@dixie-net.com(22740)

A bicyclist rode into the country for 5h. In returning, her speed was 5mi/h faster and the trip took 4h. what was her speed each way?
:
Let s = speed into the country
Then
(s+5) = speed escaping from the country
:
We know the distance out and back are the same. Write a distance equation from this fact.
Dist = time * speed
:
5s = 4(s+5)
5s - 4s + 20
s = 20 mph into the country
and
20 + 5 = 25 mph escaping from the country
:
:
Check the solutions by finding if the distances are, indeed, equal:
5(20) = 100
4(25) = 100 confirms our solutions
:
Did this make logical sense to you?